# Simultaneous Drawing of Layered Trees

### 12 Jan 2024 v3

@@ 3 Two Trees on Arbitrarily Many Layers
We prove that the vertices of $T_{2}$ are ordered according to $<^{2}_{j}$ by induction on the layer $j$, starting with the layer of the root $r$ of $T_{2}$. On layer $j=L(r_{2})$, there is only $r$, which, of course, cannot contradict $<^{2}_{j}$.
Let $j<L(r_{2})$ and $v_{1},v_{2},\dots v_{n_{2|j+1}}$ be the vertices on layer $j+1$. Then, by Claims 3 and 4, the children $C_{v_{i}}$ on layer $j$ have increasing positions respecting $<^{2}_{j}$ for every $i\in\{1,\dots,n_{2|j+1}\}$ and the edges to these children do not cross. It remains to show that no pair of edges between layers $j$ and $j+1$ without a common endpoint cross. By our induction hypothesis, for two vertices $v_{i},v_{i^{\prime}}$ on layer $j+1$ with $i<i^{\prime}$, the position $p_{i}$ of $v_{i}$ is not greater than the position $p_{i^{\prime}}$ of $v_{i^{\prime}}$. By Claim 1, it follows for the ideal positions of $p_{i}$ and $p_{i^{\prime}}$ that $p^{\star}_{i}\leq p^{\star}_{i^{\prime}}$. By Claim 2, the $\min$ and $\max$ values of the $\operatorname{P_{\textsf{opt}}}$ of the vertices on layer $j$ are monotonically increasing. Hence, by Claim 4 and the definition of $\operatorname{p_{\textsf{nat}}}$, there are no crossings between edges with target $v_{i}$ and edges with target $v_{i^{\prime}}$, as this would contradict $p^{\star}_{i}\leq p^{\star}_{i^{\prime}}$. Hence, the edges between layer $j$ and $j+1$ are planar, which concludes the induction step. ∎
@@
We show by induction over the layers $j=2,3,\dots$ that for a vertex $v$ and a position $p$, $o[v,p]$ is the minimum number of crossings induced by ($T_{1}$ and) the subtree of $T_{2}$ rooted at $v$, which we call $T_{v}$, across all drawings of $T_{v}$ when we place $v$ at position $p$. For $j=2$, this is clear as we just sum up the number of crossings induced by the edges to the leaves. Let $j\geq 3$. By Claim 4, we know that the dynamic program has selected the positions $\operatorname{p_{\textsf{nat}}}(c_{1},p)\leq\ldots\leq\operatorname{p_{\textsf{nat}}}(c_{|C_{v}|},p)$ for the children $c_{1},\dots,c_{|C_{v}|}$ of $v$. By our induction hypothesis, we know that, for each $c\in C_{v}$, $o[c,\operatorname{p_{\textsf{nat}}}(c,p)]$ corresponds to a drawing of $T_{c}$ at position $\operatorname{p_{\textsf{nat}}}(c,p)$ with the minimum number of crossings. We add up the number of crossings between layer $j-1$ and $j$, and by the formulation of the dynamic program, we know that this is again minimum across all positions of $c$. ∎
@@ 4 Multiple Trees on Three Layers
As in Sec. 3, we use the notion of positions and crossing functions, however, we slightly adjust their definitions to better suit the setting of this section. Let $\sigma$ be a permutation of $V_{2}(\mathcal{F})$ indexed by $1,2,\dots$ and respecting the partial order $\prec_{2}$. For $i\in\{1,\dots,k\}$ and some vertex $v\in V_{2}(\mathcal{F})\setminus V_{2}(T_{i})$, we denote the position (starting at 0) of $v$ within the subsequence of $\sigma$ consisting of the vertices $V_{2}(T_{i})\cup\{v\}$ by $p_{i}^{v}$. [^2] This is a generalization of the positions introduced in Sec. 3 where all positions were relative to (the given embedding of) $T_{1}$. Note that, in a drawing using $\sigma$ as $<_{2}$, we can charge every crossing to precisely two vertices of $V_{2}(\mathcal{F})$ as any crossing occurs between two edges that have two distinct endpoints on layer $2$. Now observe that for a vertex $v\in V_{2}(T_{j})$, where $j\in\{1,\dots,k\}$, the number of crossings charged to $v$ with respect to $\sigma$ depends only on $p_{i}^{v}$ for each $i\in\{1,\dots,k\}\setminus\{j\}$. Therefore, we introduce the crossing function $\operatorname{\chi}^{v}_{i}(p)$ returning the resulting number of crossings when we insert $v$ at a position $p\in\{1,\dots,n_{i|2}\}$ into the planar embedding of $T_{i}$. The number $\operatorname{\chi}_{\sigma}(v)$ of crossings charged to $v$ when using permutation $\sigma$ is then $\operatorname{\chi}_{\sigma}(v)=\sum_{i\in\{1,\dots,k\}\setminus\{j\}}\operatorname{\chi}^{v}_{i}(p_{i}^{v})$ and the total number $\operatorname{\chi}(\sigma)$ of crossings when using permutation $\sigma$ is then $\operatorname{\chi}(\sigma)=\sum_{v\in V_{2}(\mathcal{F})}\operatorname{\chi}_{\sigma}(v)/2$.
@@
Clearly, any $st$-path $\pi$ in $H$ has (unweighted) length $n_{2}$. If we traverse $\pi$, we can think of layer $2$ as being empty when we start at $s$, and then, for each edge of $\pi$, we take the corresponding vertex of $V_{2}$ and add it to layer $2$. Since edge weights equal the number of crossings the corresponding vertices would induce in this situation, finding a lightest $st$-path in $H$ means finding a crossing-minimal total order of layer $2$ (see Fig. 3c). By constructing $H$ (using Lemma 4 to compute the edge weights) and searching for an $st$-path of minimum weight, we obtain an XP-algorithm in $k$; see Theorem 4.1, which we formally prove next.

### 28 Feb 2024 v4

@@ 3 Two Trees on Arbitrarily Many Layers
We prove that the vertices of $T_{2}$ are ordered according to $<^{2}_{j}$ by induction on layer $j$, starting with the layer of the root $r$ of $T_{2}$. On layer $j=L(r_{2})$, there is only $r$, which, of course, cannot contradict $<^{2}_{j}$.
Let $j<L(r_{2})$ and $v_{1},v_{2},\dots v_{n_{2|j+1}}$ be the vertices on layer $j+1$. Then, by Claims 3 and 4, the children $C_{v_{i}}$ on layer $j$ have increasing positions respecting $<^{2}_{j}$ for every $i\in\{1,\dots,n_{2|j+1}\}$ and the edges to these children do not cross. It remains to show that no pair of edges between layers $j$ and $j+1$ without a common endpoint cross. By our induction hypothesis, for two vertices $v_{i},v_{i^{\prime}}$ on layer $j+1$ with $i<i^{\prime}$, the position $p_{i}$ of $v_{i}$ is not greater than the position $p_{i^{\prime}}$ of $v_{i^{\prime}}$. By Claim 1, it follows for the ideal positions of $p_{i}$ and $p_{i^{\prime}}$ that $p^{\star}_{i}\leq p^{\star}_{i^{\prime}}$. By Claim 2, the $\min$ and $\max$ values of $\operatorname{P_{\mathsf{opt}}}$ of the vertices on layer $j$ are monotonically increasing. Hence, by Claim 4 and the definition of $\operatorname{p_{\mathsf{nat}}}$, there are no crossings between edges with target $v_{i}$ and edges with target $v_{i^{\prime}}$, as this would contradict $p^{\star}_{i}\leq p^{\star}_{i^{\prime}}$. Hence, the edges between layer $j$ and $j+1$ are planar, which concludes the induction step. ∎
@@
We show by induction over the layers $j=2,3,\dots$ that for a vertex $v$ and a position $p$, $o[v,p]$ is the minimum number of crossings induced by ($T_{1}$ and) the subtree of $T_{2}$ rooted at $v$, which we call $T_{v}$, across all drawings of $T_{v}$ when we place $v$ at position $p$. For $j=2$, this is clear as we just sum up the number of crossings induced by the edges to the leaves. Let $j\geq 3$. By Claim 4, we know that the dynamic program has selected the positions $\operatorname{p_{\mathsf{nat}}}(c_{1},p)\leq\ldots\leq\operatorname{p_{\mathsf{nat}}}(c_{|C_{v}|},p)$ for the children $c_{1},\dots,c_{|C_{v}|}$ of $v$. By our induction hypothesis, we know that, for each $c\in C_{v}$, $o[c,\operatorname{p_{\mathsf{nat}}}(c,p)]$ corresponds to a drawing of $T_{c}$ at position $\operatorname{p_{\mathsf{nat}}}(c,p)$ with the minimum number of crossings. We add up the number of crossings between layer $j-1$ and $j$, and by the formulation of the dynamic program, we know that this is again minimum across all positions of $c$. ∎
@@ 4 Multiple Trees on Three Layers
As in Sec. 3, we use the notion of positions and crossing functions, however, we slightly adjust their definitions to better suit the setting of this section. Let $\sigma$ be a permutation of $V_{2}(\mathcal{F})$ indexed by $1,2,\dots$ and respecting the partial order $\prec_{2}$. For $i\in\{1,\dots,k\}$ and some vertex $v\in V_{2}(\mathcal{F})\setminus V_{2}(T_{i})$, we denote the position (starting at 0) of $v$ within the subsequence of $\sigma$ consisting of the vertices $V_{2}(T_{i})\cup\{v\}$ by $p_{i}^{v}$. [^2] This is a generalization of the positions introduced in Sec. 3 where all positions were relative to (the given embedding of) $T_{1}$. Note that, in a drawing using $\sigma$ as $<_{2}$, we can charge every crossing to precisely two vertices of $V_{2}(\mathcal{F})$ as any crossing occurs between two edges that have two distinct endpoints on layer $2$. Now observe that for a vertex $v\in V_{2}(T_{j})$, where $j\in\{1,\dots,k\}$, the number of crossings charged to $v$ with respect to $\sigma$ depends only on $p_{i}^{v}$ for each $i\in\{1,\dots,k\}\setminus\{j\}$. Therefore, we introduce the crossing function $\operatorname{\chi}^{v}_{i}(p)$ returning the resulting number of crossings when we insert $v$ at a position $p\in\{0,\dots,n_{i|2}\}$ into the planar embedding of $T_{i}$. The number $\operatorname{\chi}_{\sigma}(v)$ of crossings charged to $v$ when using permutation $\sigma$ is then $\operatorname{\chi}_{\sigma}(v)=\sum_{i\in\{1,\dots,k\}\setminus\{j\}}\operatorname{\chi}^{v}_{i}(p_{i}^{v})$ and the total number $\operatorname{\chi}(\sigma)$ of crossings when using permutation $\sigma$ is then $\operatorname{\chi}(\sigma)=\sum_{v\in V_{2}(\mathcal{F})}\operatorname{\chi}_{\sigma}(v)/2$.
@@
Clearly, any $st$-path $\pi$ in $H$ has (unweighted) length $n_{2}$. If we traverse $\pi$, we can think of layer $2$ as being empty when we start at $s$, and then, for each edge of $\pi$, we take the corresponding vertex of $V_{2}$ and add it to layer $2$. Since edge weights equal the number of crossings the corresponding vertices would induce in this situation, finding a lightest $st$-path in $H$ means finding a crossing-minimal total order of layer $2$ (see Fig. 3). By constructing $H$ (using Lemma 4 to compute the edge weights) and searching for an $st$-path of minimum weight, we obtain an XP-algorithm in $k$; see Theorem 4.1, which we formally prove next.